[ James is 6 years older than his sister Gessi. Two years ago, James was twice as old as Gessi. how old is James? ]
Answer by TimothyLamb(4379)

j = g + 6
j - 2 = 2(g - 2)
j - 2 = 2g - 4
---
j - 2 = 2g - 4
g + 6 - 2 = 2g - 4
---
g = 8
j = 14
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php